OpenStudy (anonymous):

simplify if possible: 2/(x-4) + 3/(x+4)

OpenStudy (anonymous):

= (2x+8+3x-12)/(x-4)(x_4) =(2x+8+3x-12)/(x^2-16) = (5x-4)/(x^2-16)

jhonyy9 (jhonyy9):

(2x+8+3x-12)/(x-4)(x+4)

OpenStudy (anonymous):

thanks!! now can you help me understand what i need to fill in the blank in this math problem

OpenStudy (anonymous):

2x-3 over x^2+2x+1= "Blank" over (x+1)^2 (x+2)

OpenStudy (anonymous):

the directions ask to write the rational expression as an equivalent expression whose denominator is the given polynomial